OpenStudy (anonymous):

Nadia has an average score of 70 after 5 tests. What is the minimum number of tests that Nadia has to score a 100 on to get the average score to at least 89?

OpenStudy (watchmath):

Let \(n\) be the number needed. So we want to have \(5\cdot70+100n\geq 89(n+5)\) \(350+100n \geq 89n+445\) \(11n\geq 95\) \(n\geq 95/11\) So at least \(n=9\).